OpenStudy (anonymous):

I need help with understanding how to solve this 3(2z+8)=-6

OpenStudy (chriss):

\[2z+8=\frac{-6}{3}\] 2z+8=-2 2z=-10 z=-5

OpenStudy (anonymous):